# Swampland Statistics for Black Holes

### v2

@@ Abstract
Keywords: Black holes, Swampland conjectures, Fermi ball, Hawking evaporation, Species scale.
@@ 1 Introduction
Recently, a special interest has been devoted to the study of effective field theories (EFTs) from higher dimensional supergravity scenarios such as superstrings and M-theory inspired models in connection with the swampland program [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The latter aims to address such EFTs through discussions on the superstring landscape. Moreover, it could be used to check for consistencies with quantum gravity (QG). Moreover, this program could provide an alternative falsification mechanism in the light of the absence of the empirical evidence for certain physical theories. A close inspection shows that many swampland criteria have been suggested, developed, and refined in the context of superstring compactifications. Generally, these compactifications generate various scalar fields in the associated spectrum. Some of them correspond to appropriate geometric concepts including the size and the shape of internal compact spaces with non-trivial holonomy groups as well as the Calabi-Yau manifolds [11]. These scenarios have been largely elaborated in order to check and verify the application of the suggested swampland conditions. In contact with such developments, two relevant conjectures have been investigated being known by distance conjecture (DC) and weak gravity conjecture (WGC). The first one concerns the implication of the stringy moduli space in the EFT building models. In superstring theory, for instance, this space is parametrized by massless scalar fields originated from the compactification mechanism dealing with the dimension reduction of the space-time in which the supergravity spectrums live. This physical road could predict certain light particles and towers at large moduli space distances according to T-duality arguments and stringy size behaviors [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. The second conjecture (WGC) was first proposed in [25]. After that, it has been developed and approached from different angles[26, 27, 28, 29, 30, 31, 32]. This conjecture stems from the fact that gravity is the weakest fundamental interaction which is translated into two main observations. First, a cutoff scale at which the low energy EFT breaks is bound from the top by a gauge coupling. Secondly, the stable black hole horizons must have little to no charge compared to black ole mass. Otherwise at extremal limits the horizon become unstable, effectively suppressing naked singularities. Moreover, the correspondence between the WGC and the DC had been explained in [33, 34, 35, 36]. Specifically, the saturation of WGC bounds by extending the moduli distance is a common observation.
@@
The aim of this work is to contribute to these activities by approaching certain black hole issues, including remnants, by providing a statistical description relaying on WGC in the swampland program. Inspired by the Pauli exclusion principal in the context of the Fermi sphere, we derive an inequality which can be utilized to verify the instability manifestation of the black holes by help of a statistical characteristic function. For a generic number of species, we show that this function is in agreement with the weak gravity swampland conjecture. Next, we examine the possibility of using the cutoff as an interval estimation problem by imposing an upper limit on the black hole mass scale, which is compatible with some results obtained in the literature. [41, 42, 43, 44, 45, 46, 47, 48]. Using the developed formalism for the proposed instability scenarios, we provide a remnant production rate suppression mechanism. Moreover, we reconsider the stability behaviors of the Reissner–Nordström black holes. Precisely, we show that the proposed instabilities prohibit naked singularity behaviors.
@@ 2 Remnant production rate
In this section, we would like to address issues associated with the contemporary theory of black holes. In particular, we discuss the corresponding remnant problems. As the name hints, the remnants are evaporating black hole remains. They have been proposed to overcome the information loss problem in black hole physics. Various attempts have been made to solve this problem. Concretely, a black hole formation description was first proposed in [49] by establishing the expression form $\frac{M_{p}^{3}}{\pi^{3}}e^{-\frac{M_{p}^{2}}{16\pi T^{2}}}$, where $M_{p}$ is the Planck mass. Later, this work has been developed by providing a classical description of such a formation[50]. In these contributions, the formation rate of the black holes has been derived, at a finite temperature. According to the works developed in [51], these scenarios have been linked to the remnant production rate via the following form
@@
where $N_{R}$ is the number of different remnant species. $M$ is the mass of the black hole. The main issue of this description is that the statistics of such objects do not add up. In fact, it has been observed that the production probability is very large for remnants at small mass scales including the Planck one. Alternatively, in order to not modify the Bekenstien-Hawking entropy, it has been argued that the remnant population could take the following form
@@
where $r$ is the distance of a fiducial observer in the thermal atmosphere [52]. In this way, the remnants are taken to be in the Planck scale. However, other mass values could be considered. Additionally, at such small scales, the distance has been identified with the internal entropy of the black hole $S_{R}$ through the relation $r=\frac{S_{R}}{M_{p}}$. Therefore, the remnant production rate can be governed by the black hole internal entropy, being an increasing quantity, muddying the water further.
@@ 3 WGC from the asymmetric instability
Secondly, it could solve the naked singularity problems by predicting the decay of the extremal black hole into EFT stable particles. [25]. In gauge theories weakly coupled to gravity, there exist charged states of a mass $m$ and an electric charge $q$ satisfying the inequality
@@
In the present investigation, we attempt to approach the problem of remnants via WGC using a statistical description. To do so, we first reconsider the study of such a conjecture through a suggested asymmetric instability scenario. This discussion will be based on the following proposed statement. In a stable (non decaying) cavity, the number of the accessible charged states should not exceed the available number of massive states representing the entirety of the system. This statement has been inspired by the Pauli exclusion principal which has been exploited to introduce the asymmetric instability (or asymmetric energy) in the semi-empirical mass model [54], where the higher energy levels are occupied by less stable states. Motivated by the Fermi description of the energy of a ball of non relativistic $\frac{1}{2}$-spin fermions (Fermi ball) [55], we would like to provide a possible description of such instabilities. Concretely, this will be done using a Fermi spherical description. In this way, the number of charged states can be expressed as
@@
Using the expressions $dx_{1}dx_{2}=r^{2}\sin(\theta)d\theta d\phi=dA$ and $dp_{1}dp_{2}=dp_{\theta}dp_{\phi}=d^{2}p$ with the condition $p_{\theta}=p_{\phi}=p$, this gives arise the following Fermi sphere law
@@
In the black hole physics, for instance, the quantity $A$ can be identified with the event horizon area of a Schwarzschild black hole, given by $A=\frac{16\pi G^{2}M^{2}}{c^{4}}$ [53]. Combining Eq.(3.5), and the relation $p=\sqrt{2m_{g}E_{z}}$, the energy associated with the asymmetric instability is found to be
@@
Using the Planck units and the global charge, we have
@@ 3.1 UV divergence and cutoffs
In this way, the mass can be expressed in terms of the standard deviation $\sqrt{Var[\frac{z\tau}{M^{2}}]}$ via the mean $\langle\frac{z\tau}{M^{2}}\rangle$ as follows
$M$ $=$ $\frac{\sqrt{z\tau}}{\sqrt{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}}=\frac{z\sqrt{\tau}}{\sqrt{2+n\sqrt{2}}}$ (3.22)
where $n$ is an integer describing the deviation degree. Accordingly, the amplitude of accessing a mass value in the rang $0,\frac{z\tau}{M^{2}}]$ is $\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$ 111 via the computation $\frac{z^{2}}{\Gamma(2)}\int_{0}^{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}Xe^{-zX}dX=\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$. where $\gamma(\alpha,x)$ is the lower incomplete gamma function given by the integral $\int^{x}_{0}x^{\alpha-1}e^{-x}dx$ [56]. Taking $n=20$, the amplitude reaches the value of $99.999\hskip 2.0pt999\hskip 2.0pt9997797\%$ for $M>\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$.Translating this for normal distributions, we conclude that the states bellow a cutoff $M_{0}=\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$ are inaccessible, with a $7\sigma$ confidence.
@@
$\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T(2+n\sqrt{2})}}(\sum\limits_{i}z_{i})<M_{0}.$ (3.23)
One can choose $N_{sp}$ to be large enough so that the term $2+n\sqrt{2}$ is ignored. According to [41], this implies that one has $\Lambda_{sp}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{0}$, where $\Lambda_{sp}$ represents the species scale. This observation is consistent with the inverse relation between the species scale $\Lambda_{sp}$ and the number of species $N_{sp}$, which could be linked to the gauge coupling via $N_{sp}\sim\frac{1}{g^{2}}$, restoring the condition given in Eq.(3.1). However, a correction term $\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})$ has appeared.
@@
This shows that WGC is consistent with the aforementioned assumptions. In connection with the previous arguments about cutoffs, the probability $\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})$ decreases by sending $M_{0}$ to $\frac{Q\sqrt{\tau}}{g}$. This can be supported by the relation
$\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})=Erf(\frac{Q\sqrt{\tau}}{gM_{0}})-Erf(1),$ (3.27)
@@ 3.2 Remnant production rate and WGC
The suppression of accessible states could be interpreted for small scale regimes. As the black hole evaporates down to a small scale, the horizon size decreases. Due to the Pauli exclusion principal, the pair produced charges can no longer exist on the black hole surface. However, the existence of such objects can be supported by a disintegrating scenario to other stable black holes and particles matching with EFTs in question. This should follow possible decay paths with different probabilities. For a black hole evaporating down to a Planck scale, the only disintegration path is to decay into stable particles as stated by WGC [25, 36].
@@ 5 Conclusion
In this paper, we have approached certain black hole issues including remnant problems by providing a statistical description based on WGC in the swampland program. It has been suggested that this conjecture basically postulates that, at small scales, the black holes should be unstable. In fact, there are possible decay channels induced by set instabilities. Based on such arguments, we have carried out a statistical investigation for non-supersymmetric four dimensional black holes, where such instabilities may be underlined by the Pauli exclusion principal. Precisely, we have shown that this statistical framework allows one to derive the charge to the mass ratio inequality $\frac{Q^{2}}{M^{2}}>\frac{32\pi g^{2}m_{g}T}{T_{p}M^{3}_{p}}$, where $m_{g}$ is the mass of a particle with a gauge coupling $g$ and $T$ is a thermodynamical equilibrium temperature with Planck scales. Consequently, it has been remarked that this formulation is similar to that of the weak gravity swampland conjecture, as the proposed instabilities also appear at small scales. Supported by the proposed instabilities, we have provided a suppression mechanism to the remnant production rate. Moreover, we have considered a generic situation to address the species scale. Specifically, we have treated the cuts in the mass scale of the black hole $M_{0}$ as an interval estimation problem. For a generic situation, this has resulted in a black hole mass cutoff lower bound $\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{0}$, where $z_{i}$ and $N_{sp}$ are the number of charged states and the number species, respectively. We have observed that the upper bound can be considered as a corrected form of the species scale introduced in the literature. Using the developed formalism of the proposed instabilities, we have reconsidered the study of the Reissner–Nordström black holes as a possible application for charged solutions. As a result, we have found that this illustrating model is consistent with the previous observations, including the remnant production rate suppression. Furthermore, we have revealed that the proposed instabilities prohibit naked singularity behaviors due to the constraint $\frac{g^{2}}{\tau}\leq G\pi\epsilon_{0}$.

### v3

@@ Abstract
Keywords: Black holes, Swampland conjectures, Fermi sphere, Hawking evaporation, Species scale.
@@ 1 Introduction
Recently, a special interest has been devoted to the study of effective field theories (EFTs) from higher dimensional supergravity scenarios such as superstrings and M-theory inspired models in connection with the swampland program [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The latter aims to address such EFTs through discussions on the superstring landscape. Moreover, it could be used to check for consistencies with quantum gravity (QG). Moreover, this program could provide an alternative falsification mechanism in the light of the absence of the empirical evidence for certain physical theories. A close inspection shows that many swampland criteria have been suggested, developed, and refined in the context of superstring compactifications. Generally, these compactifications generate various scalar fields in the associated spectrum. Some of them correspond to appropriate geometric concepts including the size and the shape of internal compact spaces with non-trivial holonomy groups as well as the Calabi-Yau manifolds [11, 12, 13, 14]. These scenarios have been largely elaborated in order to check and verify the application of the suggested swampland conditions. In contact with such developments, two relevant conjectures have been investigated being known by distance conjecture (DC) and weak gravity conjecture (WGC). The first one concerns the implication of the stringy moduli space in the EFT building models. In superstring theory, for instance, this space is parametrized by massless scalar fields originated from the compactification mechanism dealing with the dimension reduction of the space-time in which the supergravity spectrums live. This physical road could predict certain light particles and towers at large moduli space distances according to T-duality arguments and stringy size behaviors [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. The second conjecture (WGC) was first proposed in [25]. After that, it has been developed and approached from different angles[26, 27, 28, 29, 30, 31, 32]. This conjecture stems from the fact that gravity is the weakest fundamental interaction which is translated into two main observations. First, a cutoff scale at which the low energy EFT breaks is bound from the top by a gauge coupling. Secondly, the stable black hole horizons must have little to no charge compared to black ole mass. Otherwise at extremal limits the horizon become unstable, effectively suppressing naked singularities. Moreover, the correspondence between the WGC and the DC had been explained in [33, 34, 35, 36]. Specifically, the saturation of WGC bounds by extending the moduli distance is a common observation.
@@
The aim of this work is to contribute to these activities by approaching certain black hole issues, including remnants, by providing a statistical description relaying on WGC in the swampland program. Inspired by the Pauli exclusion principle in the context of the Fermi sphere, we derive an inequality which can be utilized to verify the instability manifestation of the black holes by help of a statistical characteristic function. For a generic number of species, we show that this function is in agreement with the weak gravity swampland conjecture. Next, we examine the possibility of using the cutoff as an interval estimation problem by imposing an upper limit on the black hole mass scale, which is compatible with some results obtained in the literature. [41, 42, 43, 44, 45, 46, 47, 48]. Using the developed formalism for the proposed instability scenarios, we provide a remnant production rate suppression mechanism. Moreover, we reconsider the stability behaviors of the Reissner–Nordström black holes. Precisely, we show that the proposed instabilities prohibit naked singularity behaviors.
@@ 2 Remnant production rate
In this section, we would like to address issues associated with the contemporary theory of black holes. In particular, we discuss the corresponding remnant problems. As the name hints, the remnants are evaporating black hole remains. They have been proposed to overcome the information loss problem in black hole physics. Various attempts have been made to solve this problem. Concretely, a black hole formation description was first proposed in [49] by establishing the expression form $\frac{M_{p}^{3}}{\pi^{3}}e^{-\frac{M_{p}^{2}}{16\pi T^{2}}}$, where $M_{p}$ is the Planck mass. Later, this work has been developed by providing a classical description of such a formation[50]. In these contributions, the formation rate of the black holes has been derived, at a finite temperature being related to the black hole mass via the Hawking radiation, $T=\frac{M_{p}^{2}}{8\pi M}$. According to the works developed in [51], these scenarios have been linked to the remnant production rate via the following form
@@
where $N_{R}$ is the number of different remnant species. $M$ is the mass of the black hole. The main issue of this description is that the statistics of such objects do not add up. In fact, it has been observed that the production probability is very large for remnants at small mass scales including the Planck one. In any case, the number of remaining species can be infinite, as there is no fixed limit for this number. In certain cases, it has been observed that the situation can be exacerbated when the internal entropy of the remains is used to count $N_{R}$. In order to avoid altering the Bekenstien-Hawking entropy, it has been proposed that the remaining population could take the following form
@@
where $r$ is the distance of a fiducial observer in the thermal atmosphere [52]. In this way, the remnants are taken to be in the Planck scale. However, lower mass values could be considered. Additionally, at such small scales, the distance has been identified with the internal entropy of the black hole $S_{R}$ through the relation $r=\frac{S_{R}}{M_{p}}$. Therefore, the remnant production rate can be governed by the black hole internal entropy, being an increasing quantity, muddying the water further.
From another perspective, in [53], it was pointed out that when the content of a black hole is separated into several quantum states, called “informons,” the decay rate of a weakly coupled electric field into these informons (the rate of residue production) includes a divergent term in the form of a sum of exponential expressions over the number of these states. This divergence would be caused by the infinite number of informons whose mass is less than the Planck scale. As a prelude to what follows, this question will be addressed by proposing that it is not possible to fit an infinite number of states into the black hole. Alternatively, the usual Hawking evaporation would be interrupted by the decay of the black hole along certain probable decay channels.
@@ 3 WGC from the asymmetric instability
Secondly, it could solve the naked singularity problems by predicting the decay of the extremal black hole into EFT stable particles. [25]. Explicitly, in gauge theories weakly coupled to gravity, there exist charged states of a mass $m$ and an electric charge $q$ satisfying the inequality
@@
In the present investigation, we attempt to approach the problem of remnants via WGC using a statistical description. To do so, we first reconsider the study of such a conjecture through a suggested asymmetric instability scenario. This discussion will be based on the following proposed statement. In a stable (non decaying) cavity, the number of the accessible charged states should not exceed the available number of massive states representing the entirety of the system. This statement has been inspired by the Pauli exclusion principle which has been exploited to introduce the asymmetric instability (or asymmetric energy) in the semi-empirical mass model [54], where the higher energy levels are occupied by less stable states. Motivated by the Fermi description of the energy of a ball of non relativistic $\frac{1}{2}$-spin fermions [55], we would like to provide a possible description of such instabilities. Concretely, this will be done using a Fermi spherical description. In this way, the number of charged states can be expressed as
@@
In the following, we use the expressions $dx_{1}dx_{2}=r^{2}\sin(\theta)d\theta d\phi=dA$ and $dp_{1}dp_{2}=dp_{\theta}dp_{\phi}=d^{2}p$ with the condition $p_{\theta}=p_{\phi}=p$. By deriving the number of states with respect to the momentum and integrating over the spatial variables, we arrive at the following Fermi sphere law
@@
This law basically dictate how one should count the states in the given phase space.
In the black hole physics, for instance, the quantity $A$ can be identified with the event horizon area of a Schwarzschild black hole, given by $A=\frac{16\pi G^{2}M^{2}}{c^{4}}$ [56]. Since we wish to establish a scenario of instability governed by a charge-to-mass ratio, we will not delve into more detailed phenomenological studies of these states. We will therefore consider that the observer counting the states is in a reference frame where relativistic effects are negligible. Combining Eq.(3.5), and the relation $p=\sqrt{2m_{g}E_{z}}$, the energy associated with the asymmetric instability is found to be
@@
Considering quantized total charge $Q=zg$, and using the Planck units and the global charge, we have
@@ 3.1 UV divergence and cutoffs
In this way, the mass can be treated as a step indexed by an integer. Thus it is expressed in terms of the standard deviation $\sqrt{Var[\frac{z\tau}{M^{2}}]}$ via the mean $\langle\frac{z\tau}{M^{2}}\rangle$ as follows
$M_{n}$ $=$ $\frac{\sqrt{z\tau}}{\sqrt{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}}=\frac{z\sqrt{\tau}}{\sqrt{2+n\sqrt{2}}}$ (3.22)
where $n$ is an integer indicating the degree of deviation from the mean. Accordingly, the amplitude of accessing a mass value in the rang $0,\frac{z\tau}{M^{2}}]$ is $\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$ 111via the computation $\frac{z^{2}}{\Gamma(2)}\int_{0}^{\frac{2+n\sqrt{2}}{z}}Xe^{-zX}dX=\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$. where $\gamma(\alpha,x)$ is the lower incomplete gamma function given by the integral $\int^{x}_{0}x^{\alpha-1}e^{-x}dx$ [57]. As a simple illustration, Fig.(1) shows the selected sample by considering a degree of deviation $n=1$. Precisely, with such a degree of deviations, one can access a state in the range $\frac{M_{1}}{\sqrt{\tau}},\infty[$ with a probability of $85.5\%$.
Figure 1: Cutoff mas $M_{1}$ at a degree $n=1$ in the $z=2$ case.
Additionally, by taking $n=20$, the amplitude reaches the value of $99.999\hskip 2.0pt999\hskip 2.0pt9997797\%$ for $M\geq\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$.Translating this for normal distributions, we conclude that the states bellow a cutoff $M_{20}=\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$ are inaccessible, with a $7\sigma$ confidence.
@@
$\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T(2+n\sqrt{2})}}(\sum\limits_{i}z_{i})<M_{n}.$ (3.23)
One can choose $N_{sp}$ to be large enough so that the term $2+n\sqrt{2}$ is ignored. According to [41], this implies that one has $\Lambda_{sp}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{n}$, where $\Lambda_{sp}$ represents the species scale. This observation is consistent with the inverse relation between the species scale $\Lambda_{sp}$ and the number of species $N_{sp}$, which could be linked to the gauge coupling via $N_{sp}\sim\frac{1}{g^{2}}$, restoring the condition given in Eq.(3.1). However, a correction term $\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})$ has appeared.
@@
This shows that WGC is consistent with the aforementioned assumptions. In connection with the previous arguments about cutoffs, the probability $\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})$ decreases by sending a cutoff $M^{\star}$ to $\frac{Q\sqrt{\tau}}{g}$. This can be supported by the relation
$\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})=erf(\frac{Q\sqrt{\tau}}{gM^{\star}})-erf(1),$ (3.27)
@@ 3.2 Remnant production rate and WGC
The suppression of accessible states could be interpreted for small scale regimes. As the black hole evaporates down to a small scale, the horizon size decreases. Due to the Pauli exclusion principle, the pair produced charges can no longer exist on the black hole surface. However, the existence of such objects can be supported by a disintegrating scenario into other stable black holes and particles matching with EFTs in question. This should follow possible decay paths with different probabilities. For a black hole evaporating down to a Planck scale, the only disintegration path is to decay into stable particles as stated by WGC [25, 36].Lastly, the density function introduced should contribute to the Bekenstein-Hawking entropy, as shown in equation (3.28). This leads to significant results concerning the Bekenstein bound, where states are counted over a cubical box. However, the implications of this work for entropy will be the subject of future investigations.
@@ 5 Conclusion
In this paper, we have approached certain black hole issues including remnant problems by providing a statistical description based on WGC in the swampland program. It has been suggested that this conjecture basically postulates that, at small scales, the black holes should be unstable. In fact, there are possible decay channels induced by set instabilities. Based on such arguments, we have carried out a statistical investigation for non-supersymmetric four dimensional black holes, where such instabilities may be underlined by the Pauli exclusion principle. Precisely, we have shown that this statistical framework allows one to derive the charge to the mass ratio inequality $\frac{Q^{2}}{M^{2}}>\frac{32\pi g^{2}m_{g}T}{T_{p}M^{3}_{p}}$, where $m_{g}$ is the mass of a particle with a gauge coupling $g$ and $T$ is a thermodynamical equilibrium temperature with Planck scales. Consequently, it has been remarked that this formulation is similar to that of the weak gravity swampland conjecture, as the proposed instabilities also appear at small scales. Supported by the proposed instabilities, we have provided a suppression mechanism to the remnant production rate. Moreover, we have considered a generic situation to address the species scale. Specifically, we have treated the cuts in the mass scale of the black hole $M_{0}$ as an interval estimation problem. For a generic situation, this has resulted in a black hole mass cutoff lower bound $\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{n}$, where $z_{i}$ and $N_{sp}$ are the number of charged states and the number species, respectively. We have observed that the upper bound can be considered as a corrected form of the species scale introduced in the literature. Using the developed formalism of the proposed instabilities, we have reconsidered the study of the Reissner–Nordström black holes as a possible application for charged solutions. As a result, we have found that this illustrating model is consistent with the previous observations, including the remnant production rate suppression. Furthermore, we have revealed that the proposed instabilities prohibit naked singularity behaviors due to the constraint $\frac{g^{2}}{\tau}<G\pi\epsilon_{0}$.
@@
Conflicts of interest
The authors declare that they have no conflicts of interest.